# Probing Light Bosonic Dark Matter with Transmon Qubits

### v1

@@ IV Discussion
Ultralight bosonic particles have emerged as compelling candidates for dark matter (DM) due to their potential to address various cosmological and astrophysical mysteries. Detecting these elusive particles, however, presents a significant challenge to the scientific community. Recent studies and technological advancements are pushing the boundaries of what is possible. The use of quantum technologies in various cavity experiments, antenna experiments and collective excitation experiments, etc., and the exploration of novel detection methods are paving the way for potentially groundbreaking discoveries in our understanding of DM. Recent study showed that quantum qubit, a foundation for exploring quantum computing, can be made into competitive devices for direct detections of light DM. In this paper, we explored the constraints on the absorbable DM from the measurement of quasiparticles in the transmon qubit, following the strategy of Ref. Das et al. (2024b). Our result shows the first constraints on these DMs from transmon qubit. These developments underscore the growing interest and progress in using quantum technologies, particularly transmon qubits, for the detection of dark matter. As research continues, we can expect further refinements in the sensitivity and capabilities of these devices, potentially leading to groundbreaking discoveries in our understanding of dark matter.
In prospect, developments in quantum qubits can provide a promising avenue for light bosonic dark matter direct detection with more dedicated R&D. In this paper the projective limit is a proof-of-principle estimate based on the sensitivity to the equilibrium density of quasi-particles in aluminum. Meanwhile, new detectors utilizing the fundamental architecture of transmon qubits are quickly advancing toward single Cooper-pair breaking sensitivity Fink et al. (2023); Li et al. (2024); Ramanathan et al. (2024), demonstrating a notable reduction in the detection threshold and a substantial enhancement in sensitivity. When single quasi-particle sensitivity becomes a reality, direct measurement of the DM-induced production rate with a larger test mass and over longer exposure periods will likely bring significant improvement in the detection sensitivity.

### v2

@@ IV Discussion
Ultralight bosonic particles have emerged as compelling candidates for dark matter (DM) due to their potential to address various cosmological and astrophysical mysteries. Detecting these elusive particles, however, presents a significant challenge to the scientific community. Recent studies and technological advancements are pushing the boundaries of what is possible. The use of quantum technologies in various cavity experiments, antenna experiments and collective excitation experiments, etc., and the exploration of novel detection methods are paving the way for potentially groundbreaking discoveries in our understanding of DM. Recent studies showed that quantum qubit, a foundation for exploring quantum computing, can serve as competitive devices for direct detection of light DM. In this paper, we explored the constraints on light bosonic DM absorption from the measurement of quasiparticles in the transmon qubit. Our result gives the first constraints on dark photon, scalar and axion-like particle dark matter from existing transmon qubit measurements. This underscores the growing interest and progress in quantum technologies to play a role in the search for dark matter.
In prospect, developments in quantum qubits can provide a promising avenue for light bosonic dark matter direct detection with more dedicated R&D. In this paper the projective limit is a proof-of-principle estimate based on the sensitivity to the equilibrium density of quasi-particles in aluminum. Meanwhile, new detectors utilizing the fundamental architecture of transmon qubits are quickly advancing toward single Cooper-pair breaking sensitivity Fink et al. (2023); Li et al. (2024); Ramanathan et al. (2024), demonstrating a notable reduction in the detection threshold and a substantial enhancement in sensitivity. When single quasi-particle sensitivity becomes a reality, direct measurement of the DM-induced production rate with a larger test mass and over longer exposure periods will likely improve the detection sensitivity significantly.